OpenStudy (anonymous):

Suppose a computer virus begins by infecting 8 computers in the first hour after it is released. Each hour after that, each newly infected computer causes 8 more computers to become infected. The function y = 8x models this situation. Make a table with integer values of x from 1 to 4

OpenStudy (ghazi):

put x=1 you'll have y= 8*1, similarly put x=2, you'll have y= 8*2..and so on